OpenStudy (anonymous):

What is the slope of the line represented by the equation ? y= x/3 - 5 -5 3 1/3 -3/5

OpenStudy (anonymous):

y = mx + b where m is the slope. Can you take on with the rest?

OpenStudy (anonymous):

mx = (1/3)x what do you think the value of m?

OpenStudy (anonymous):

divide by x by both sides, you will get the value of m.

OpenStudy (anonymous):

give me a medal atleast.

OpenStudy (anonymous):

I don't know how to do that and I don't know how to give you a medal.

OpenStudy (anonymous):

Divide X from both sides. How do I do that?

OpenStudy (anonymous):

@Boblovesmath

OpenStudy (anonymous):

\[mx = (1/3)x\] --------> dividing my x \[m = 1/3\]

OpenStudy (anonymous):

There you go, should've just done that in the first place. Thank you.

OpenStudy (anonymous):

You want the answer without any effort. I shouldn't have done that from the first place.

OpenStudy (anonymous):

No I wanted you to show me how to do it so I can teach my cousin, you didn't have to give the answer at all. It helped me out that you did though so thanks.

OpenStudy (anonymous):

Okay, you are welcome then :D

OpenStudy (anonymous):

:D